OpenStudy (anonymous):

find the domain of the function f(x)=log2(25-x^2 step by step please...:)

OpenStudy (anonymous):

\[\log _{2}(25-x^2)=\log _{2}5^2-\log _{2}x^2\]

OpenStudy (anonymous):

\[5^2-x^2\]\[25=x^2\]\[x=\pm \sqrt{25}\]\[x= \pm 5\]